# Differentiable Modal Synthesis for Physical Modeling of Planar String Sound and Motion Simulation

### v1

@@ Abstract
While significant advancements have been made in music generation and differentiable sound synthesis within machine learning and computer audition, the simulation of instrument vibration guided by physical laws has been underexplored. To address this gap, we introduce a novel model for simulating the spatio-temporal motion of nonlinear strings, integrating modal synthesis and spectral modeling within a neural network framework. Our model leverages physical properties and fundamental frequencies as inputs, outputting string states across time and space that solve the partial differential equation characterizing the nonlinear string. Empirical evaluations demonstrate that the proposed architecture achieves superior accuracy in string motion simulation compared to existing baseline architectures. The code and demo are available online.  [^1] https://bit.ly/dmsp-notion-demo  [^2] https://huggingface.co/spaces/szin94/dmsp
@@ 1 Introduction
Despite the growing recognition of DDSP as a promising sound synthesis methodology, its extension to physical modeling remains underexplored. In the context of musical sound synthesis, some attempts have been made to synthesize rigid-body contact sounds of arbitrary shapes using neural networks supervised by finite-element method (FEM) solvers. Jin et al. [29, 30] propose a neural network that predicts contact sounds from voxelized objects, inspired by the modal technique that synthesizes sound using eigenvalues and eigenvectors. Diaz et al. [31] leverage a differentiable infinite impulse response filter to synthesize contact sounds from rasterized occupancy grids in an end-to-end manner. These methods can interactively synthesize sounds for various contact conditions and materials with notable efficiency, circumventing the need for an offline optimization process typical of modal techniques. However, these methods, which resort to the FEM solver, are vulnerable in modeling the dynamic behavior (e.g., glissando, vibrato) or in simulating the motion of the object. This vulnerability demands the need for further research to enhance the capability of neural networks in physical modeling for more dynamic and complex sound synthesis scenarios.
@@
  • To the best of our knowledge, this is the first approach that is differentiable and can synthesize the motion and the sound of a musical string with a dynamic control over the pitch and the material properties.
@@
Figure 1: System overview. The DMSP model encodes the physical properties of a string (e.g., tension, stiffness, damping, and initial conditions) to estimate the displacement of the string plucked at pitch $f_{0}$ at a given time $t\in[0,\infty)$ and position $x\in\Omega$. By concatenating the DMSP outputs over the domain $(x,t)\in\Omega\times[0,\infty)$, the simulated motion of the string can be visualized. Sampling the outputs at a particular position $x$ allows one to hear the synthesized string sound, akin to listening with a stethoscope at the pickup position.
@@ 2.1 Physical Modeling of Musical String Instrument
The linear string of its length $L$, vibrating with wave speed $\gamma$, stiffness $\kappa$, and frequency-independent damping factor $\sigma_{0}$, is described by Equation 1. Given the initial conditions (IC) for $x\in\Omega=[-L/2,+L/2]$ as $u(x,0)=u_{0}(x)$ and $\partial_{t}u(x,0)=0$, and appropriate boundary conditions (BC), the corresponding solution $u(x,t)$ represents the motion of a damped linear stiff string. Particularly, for a clamped boundary condition, i.e., $u(\pm L/2,t)=\partial_{x}u(\pm L/2,t)=0$ for all $t\in[0,\infty)$, an analytic solution can be obtained as follows.
@@
The derivation of Equation 2 can be found in Appendix C. The allowed values of $\mu_{n}$ and $\nu_{n}$ are determined by the boundary conditions, while the coefficients $c_{1}$ and $c_{2}$ are determined using the initial condition $\sum_{n=1}^{\infty}X_{n}(x)=u_{0}(x)$. Determining these values typically requires an offline numerical solving process, where the obtained values can be stored in memory for real-time computation of $u(x,t)$. The stiffness modeled by the 4th order derivative induces a hyperbolic solution, resulting in non-integer multiples of the mode frequencies $\omega_{n}$, which leads to physical inharmonicity. The damping factor $\sigma_{0}$ causes an exponential decay in the temporal amplitude.
@@
Nonlinear Damped Stiff String. The generalization of the linear wave Equation 1 to nonlinear string vibrations was first introduced by Kirchhoff [32] and Carrier [33]. The Kirchhoff–Carrier system models elastic strings in two dimensions, and when extended so that transverse and longitudinal motions are coupled, phantom partials can be exhibited, resulting in a richer timbre [34]. A model of such planar string vibration is as follows [35, 36].
@@ 2.2 Differentiable Digital Signal Processing
Table 1: Comparison between methods.
@@
Physical Nonlinear Differentiable Pre-computation Inference
@@
FDTD - $\mathcal{O}(N_{x}N_{t})$
DDSP - $\mathcal{O}(N_{r})$
@@ 3.2 Network Architecture
Parameter Encoder. To effectively capture material features inherent in the PDE parameter values, the parameter encoder leverages a random Fourier feature (RFF) layer [48, 49]. Given T60 frequencies $f_{\mathrm{T60}}^{(i)}$ and their corresponding times $t_{\mathrm{T60}}^{(i)}$ for $i=1,2$, the frequency-dependent damping coefficients $\sigma_{0}$ (and $\sigma_{1}$, if applicable) are derived using Equation 24. Subsequently, the frequency-independent damping factor $\exp(-\sigma_{0}t)$ is computed, explicitly multiplied by the mode amplitudes. All PDE parameters $\rho=\left\{\kappa,\alpha,\sigma_{0},\sigma_{1}\right\}\in\mathbb{R}^{4}\subset\mathcal{P}$ are encoded into a feature vector $h\in\mathbb{R}^{4\times d}$ with a Fourier embedding size of $d=256$.
AM and FM Blocks. As illustrated in Figure 3, DMSP employs amplitude modulation (AM) and frequency modulation (FM) modules to modulate the mode frequencies and amplitudes of the linear solution, as depicted in Equation 2, to synthesize the solution of the nonlinear wave described in Equation 3. We utilize two multilayer perceptrons (MLPs) for the modulation layers. Although a simpler and perhaps more conventional choice of architecture would be a GRU, to the decoder architecture of DDSP [16], we choose MLPs for their slightly better empirical results. It’s noteworthy that DDSP decodes the sinusoidal frequency envelope with fixed frequency values. In contrast, DMSP decodes both the envelope and the frequency values independently, employing two distinct MLP blocks, namely AM and FM.
@@ 3.3 Loss Function
In training DMSP, we employ a combination of four loss terms: (1) waveform $\ell_{1}$ loss that measures the $L_{1}$ discrepancy between the synthesized waveform and the ground truth waveform, (2) Multi-scale spectral (MSS) loss that captures spectral differences across multiple scales, ensuring fidelity in spectral representation, and (3) Pitch loss ($\mathcal{L}_{f_{0}}$) that penalizes deviations from the ground truth fundamental frequency ($f_{0}$). Optimizing the frequency parameters of sinusoidal oscillators via gradient descent over an audio loss function poses a challenge due to the non-convex nature of the optimization problem, as highlighted in various studies [50, 51]. Damped sinusoids offer a workaround for the issue of non-convexity concerning frequency parameters [51], or alternative metrics are proposed to mitigate the risk of falling into bad local minima [52, 53]. In our approach, we adopt a parameter regression joint training strategy, akin to the pre-training phase of the work by Engel et al. [54]. Among the output mode frequencies, we optimize one frequency component to match the target fundamental frequency ($f_{0}$) annotated using CREPE [55]. As the comparison models are all aware of the frequency of the mode they are outputting, the models are updated to minimize the discrepancy between the fundamental frequency estimated as the lowest one of the modes (denoted by $\hat{f}_{0}$) and the ground-truth $f_{0}$ of the FDTD-simulated audio estimated via CREPE as $\mathcal{L}_{f_{0}}=\|\hat{f}_{0}-f_{0}\|_{1}$.
@@ 4 Experiments
This section introduces the empirical investigation that seeks to find the answers to the following questions: (1) how much does the modal synthesis-based inductive bias of a neural network advantage physical modeling sound synthesis, (2) how much does it improve nonlinearity and what physical analysis is possible, and (3) to what extent are techniques required for training such model.
@@ 4.1 Experimental Setup
Baselines. As the first attempt to tackle the task of neural audio synthesis from the physical properties, we compare our proposed model to three other models. We consider three baselines, namely: Modal and DDSPish, and DDSPish without FM. Modal synthesis is the linear wave solution as in Equation 1, where the modal frequencies and the shape functions are pre-computed. DDSPish is a neural network based on the harmonic plus noise model, similar to DDSP. Yet, this -ish suffix emphasizes that this model is different from the DDSP model [16]. DDSPish does not have a reverb module but instead adds frequency modulation, and most notably it has a parameter encoder that allows generating sounds from physical parameters. Please see Appendix D for more details of the baselines.
Evaluation Metrics. We report results on three metrics, signal-distortion-ratio (SDR), scale-invariant signal-distortion-ratio (SI-SDR) [56], multi-scale spectral (MSS) distance, and the pitch difference in Hz. The MSS metric for the evaluation was computed using the short-time Fourier transformation (STFT) with 513 number of frequency bins and the hop size of 256 samples. The Pitch metric is computed as the $\ell^{1}$ norm of the difference between $\hat{f}_{0}$ and $f_{0}$.
@@ 4.2 Results
Differentiable Sound Synthesis. The efficacy of DMSP is studied as shown in Table 2. First, the modal synthesis method calculates a linear solution for a damped stiff string, and the mode for the linear test set is calculated offline. The discrepancy between the outcomes of modal synthesis and FDTD in the linear case can be attributed to the absence of frequency-dependent damping in Equation 1. In other words, the frequency information serves as an oracle for the Modal, as evidenced by its lowest pitch score, but the decay of amplitude over time is not sufficiently modeled, which is where the remaining models demonstrate superior performance. Considering the DDSPish models, the MSS is approximately 1.6 dB ahead of the DMSPs, but the difference can reach up to 44 dB in the case of the SI-SDR. It is worth noting that $\alpha$ is uniformly sampled within the range $(1,25)$ for the training data, falling into the category of the nonlinear strings. In the case of nonlinear test set, the difference between the modal solution and the FDTD solution becomes larger. On the other hand, DDSPish models based on spectral modeling synthesis but learned without using any modal information show the lowest performance. The superiority of DMSP is even more pronounced in the nonlinear case. While Modal, which can only cover the solution to the linear case, shows attenuated performance, the DMSPs demonstrate the best performance in all metrics. In particular, DMSP-Hybrid, which precomputes the mode frequency like Modal, performs FM and AM for the nonlinear solution, showing an SI-SDR improvement of nearly 31 dB and an MSS improvement of nearly 13 dB over Modal. DMSP, which estimates the mode as a neural network without the pre-computation step, also outperforms Modal on all metrics for the nonlinear case. The primary reason for the performance discrepancy between the nonlinear case and the linear case for DMSP is that the training data is rarely precisely equal to $1$ in the $\alpha$ distribution when it is sampled.
@@
Figure 5: Objective scores over physical parameters.
@@
Controllable Physical Simulation. The quantitative scores for various physical condition parameters are visualized in Figure 5. Trends show how the results of Modal synthesis and DMSP vary for different pickup positions ($x$), stiffness ($\kappa$), tension ($\alpha$), pluck amplitude ($p_{a}$), and pluck position ($p_{x}$). Of these, $\alpha$ and $p_{a}$ in particular are known to increase the nonlinearity of the string as they increase in magnitude, which can be seen by the lower Modal synthesis scores. For DMSP, we see an overall improvement in the score, with a lower propensity for nonlinearity. Figure 4 depicts the simulated state of the string as the pluck position in the initial condition is varied. The results synthesized by DMSP can reconstruct a very accurate initial condition, similar to the results simulated by FDTD. The vibration propagating through time along the string exhibits a distinct behavior contingent upon the initial condition. FDTD employs a recursive calculation of displacement, necessitating a number of iterations equal to the number of samples at the audio sampling rate. In contrast, DMSP is capable of obtaining the desired displacement in both time and space simultaneously.
@@
String Sound Synthesis. Spectrograms of the test samples are visualized in Figure 6. The instantaneous frequencies are identified in a rainbow color map, where the color intensities represent the logarithmic magnitude of the power spectra. The FDTD-simulated spectrogram clearly shows pitch glide and phantom partials at the beginning of the pluck. In contrast, modal synthesis methods that model linear solutions do not show these nonlinear characteristics. The DDSPish-XFM model employs a harmonic pitch skeleton comprising integer multiples of $f_{0}$, thereby precluding the inharmoicities resulting from stiffness. The DDSPish model demonstrates enhanced mode estimation capabilities through learned FM, which modulates the harmonic pitch skeleton to be inharmonic. However, there is scope for further improvement in frequency estimation, particularly in instances where the FM learning process is unstable for high frequencies. On the other hand, the DMSP model, which estimates the mode frequency and amplitude from $u_{0}$, shows an improved pitch skeleton and stable frequency estimation. The DMSP-Hybrid model, which learns to AM and FM the sinusoidal oscillators of the modal solution that requires mode precomputation, shows the most similar results to FDTD.
@@
Ablation Study. The ablation study on the choice of the training loss functions is presented in Table 3. Overall, for linear strings, the performance does not vary much depending on which loss is used, especially for pitch. This is due to the alpha masking applied to the FM block, which is designed to prevent FM from occurring when alpha is 1. This is the main factor that determines the nonlinearity of the string. For AM, there is no masking depending on alpha, so the remaining metrics except pitch do vary. For nonlinear data, the learning results vary depending on the loss design.
@@
Motion Synthesis. The main advantage of this neural net based on modal synthesis is that it can synthesize not only sound but also motion, which is one of the main characteristics of physical modeling techniques. In particular, the DMSP is able to visualize the corresponding string motion as a video, despite the fact that the receptive field and computational complexity required to obtain a solution $u(x,t)$ for a single spatio-temporal point is of order 1, as shown in Table 1, when these solutions are pooled for a given $x\in\Omega$ and $t\in[0,1)$, the corresponding string motion can be visualized as a video. Figure 7 visualizes the resulting transverse displacement of the string over time (horizontal axis) and space (vertical axis). As mentioned earlier, the transverse displacement of the FDTD is coupled to the longitudinal motion, which is why it differs from the Modal synthesis output, which synthesizes the motion of a linear damped stiff string. The results output by DMSP show improved accuracy.

### v2

@@ Abstract
While significant advancements have been made in music generation and differentiable sound synthesis within machine learning and computer audition, the simulation of instrument vibration guided by physical laws has been underexplored. To address this gap, we introduce a novel model for simulating the spatio-temporal motion of nonlinear strings, integrating modal synthesis and spectral modeling within a neural network framework. Our model leverages physical properties and fundamental frequencies as inputs, outputting string states across time and space that solve the partial differential equation characterizing the nonlinear string. Empirical evaluations demonstrate that the proposed architecture achieves superior accuracy in string motion simulation compared to existing baseline architectures. The code and demo are available online.  [^1] https://huggingface.co/spaces/szin94/dmsp
@@ 1 Introduction
Despite the growing recognition of DDSP as a promising sound synthesis methodology, its extension to physical modeling remains underexplored. Schlecht et al. [29] have presented physical modeling using Fourier Neural Operator (FNO). They train recurrent-type FNOs to learn state transitions from data spanning a few initial timesteps in the simulation and then test generalization to the subsequent long-range data. Although they have demonstrated encouraging outcomes, their approach has room for improvement in that it is unconditional, meaning that it is challenging to generalize over dynamic scenarios (e.g., glissando, vibrato). In the context of rigid-body contact sound synthesis [30, 31, 32], a few studies investigated the efficacy of training neural networks supervised by finite-element method (FEM) solvers. Jin et al. [33, 34] propose a neural network that predicts contact sounds from voxelized objects, inspired by the modal technique that synthesizes sound using eigenvalues and eigenvectors. Diaz et al. [35] leverage a differentiable infinite impulse response filter to synthesize contact sounds from rasterized occupancy grids in an end-to-end manner. These methods can interactively synthesize sounds for various contact conditions and materials with notable efficiency, circumventing the need for an offline optimization process typical of modal techniques. However, these methods, which resort to the FEM solver, are vulnerable in modeling the dynamic behavior or in simulating the motion of the object.
Figure 1: System overview. The DMSP model encodes the physical properties of a string (e.g., tension, stiffness, damping, and initial conditions) to estimate the displacement of the string plucked at pitch $f_{0}$ at a given time $t\in[0,\infty)$ and position $x\in\Omega$. By concatenating the DMSP outputs over the domain $(x,t)\in\Omega\times[0,\infty)$, the simulated motion of the string can be visualized. Reading the outputs at a particular position $x$ allows hearing the synthesized string sound, akin to listening with a stethoscope at the pickup position.
@@
  • To the best of our knowledge, this is the first differentiable approach that can synthesize the motion and the sound of musical strings with a dynamic control over the pitch and the material properties.
@@ 2.1 Physical Modeling of Musical String Instrument
The linear string of its length $L$, vibrating with wave speed $\gamma$, stiffness $\kappa$, and frequency-independent damping factor $\sigma_{0}$, is described by Equation 1. Given the initial conditions (IC) for $x\in\Omega=[-L/2,+L/2]$ as $u(x,0)=u_{0}(x)$ and $\partial_{t}u(x,0)=0$, and appropriate boundary conditions (BC), the corresponding solution $u(x,t)$ represents the motion of a damped linear stiff string. Particularly, for a clamped boundary condition, i.e., $u(\pm L/2,t)=\partial_{x}u(\pm L/2,t)=0$ for all $t\in[0,\infty)$, a modal solution can be obtained as follows.
@@
The derivation of Equation 2 can be found in Appendix A. The allowed values of $\mu_{n}$ and $\nu_{n}$ are determined by the boundary conditions, while the coefficients $c_{1}$ and $c_{2}$ are determined using the initial condition $\sum_{n=1}^{\infty}X_{n}(x)=u_{0}(x)$. Determining these values typically requires an offline numerical solving process, where the obtained values can be stored in memory for real-time computation of $u(x,t)$. The stiffness modeled by the 4th order derivative induces a hyperbolic solution, resulting in non-integer multiples of the mode frequencies $\omega_{n}$, which leads to physical inharmonicity. The damping factor $\sigma_{0}$ causes an exponential decay in the temporal amplitude.
@@
Nonlinear Damped Stiff String. The generalization of the linear wave Equation 1 to nonlinear string vibrations is first introduced by Kirchhoff [36] and Carrier [37]. The Kirchhoff–Carrier system models elastic strings in two dimensions, and when extended so that transverse and longitudinal motions are coupled, phantom partials can be exhibited, resulting in a richer timbre [38]. A model of such planar string vibration is as follows [39, 40].
@@ 2.2 Differentiable Digital Signal Processing
Table 1: Comparison between methods. Computational complexity refers to the inference scenario.
@@
Physical Nonlinear Differentiable Pre-computation Synthesis
@@
FDTD N/A $\mathcal{O}(N_{x}N_{t})$
DDSP N/A $\mathcal{O}(N_{r})$
@@ 3.2 Network Architecture
Parameter Encoder. To effectively capture material features inherent in the PDE parameter values, the parameter encoder leverages a random Fourier feature (RFF) layer [52, 53]. Given T60 frequencies $f_{\mathrm{T60}}^{(i)}$ and their corresponding times $t_{\mathrm{T60}}^{(i)}$ for $i=1,2$, the frequency-dependent damping coefficients $\sigma_{0}$ (and $\sigma_{1}$, if applicable) are derived using Equation 24. The frequency-independent damping factor $\exp(-\sigma_{0}t)$ is computed explicitly multiplied by the mode amplitudes. All PDE parameters $\rho=\left\{\kappa,\alpha,\sigma_{0},\sigma_{1}\right\}\in\mathbb{R}^{4}\subset\mathcal{P}$ are encoded into a feature vector $h\in\mathbb{R}^{4\times d}$ with a Fourier embedding size of $d=256$.
AM and FM Blocks. As illustrated in Figure 3, DMSP employs amplitude modulation (AM) and frequency modulation (FM) modules 222 It is worth noting that there are some subtle differences from the AM/FM techniques commonly used in sound synthesis and the usage of the terminology in this paper. Relying on terminology commonly used in AM/FM technique, AM/FM in this paper can be described analogically as a carrier sinusoid whose center frequency and amplitude are determined by modal decomposition and varied by a non-sinusoidal modulator. While typical AM/FM synthesis also uses non-sinusoidal modulators, the intent of this paper is different in that it borrows these for the purpose of limiting the undesired periodicity in those modulators. to modulate the mode frequencies and amplitudes of the linear solution, as depicted in Equation 2, to synthesize the solution of the nonlinear wave described in Equation 3. We utilize two multilayer perceptrons (MLPs) for the modulation layers. Although a simpler and perhaps more conventional choice of architecture would be a GRU, to the decoder architecture of DDSP [16], we choose MLPs for their slightly better empirical results. It’s noteworthy that DDSP decodes the sinusoidal frequency envelope with fixed frequency values. In contrast, DMSP decodes both the envelope and the frequency values independently, employing two distinct MLP blocks, namely AM and FM.
@@ 3.3 Loss Function
We employ a combination of four loss terms: (1) waveform $\ell_{1}$ loss ($\mathcal{L}_{1}$) that measures the $L_{1}$ discrepancy between the synthesized waveform and the ground truth waveform, (2) Multi-scale spectral (MSS) loss that captures spectral differences across multiple scales, ensuring fidelity in spectral representation, (3) Pitch loss ($\mathcal{L}_{f_{0}}$) that penalizes deviations from the ground truth fundamental frequency ($f_{0}$), and (4) Mode loss ($\mathcal{L}_{m}$), which measures the $L_{1}$ distance of the mode frequency and mode amplitude from the output of the mode estimator (if applicable) and the mode frequency and mode amplitude obtained via modal decomposition. MSS loss has been adopted as a metric for reconstruction in most neural net-based synthesis techniques [16, 21, 11]. Concerning that measuring MSS with magnitudes is not phase-sensitive by definition, we employ the $\mathcal{L}_{1}$ loss in the waveform to train the causality in spatio-temporal wave propagations. Challenges in optimizing the frequency parameters of sinusoidal oscillators via gradient descent over the spectral loss functions, due to the non-convex nature of the optimization problem, have been highlighted in various studies [54, 55]. Damped sinusoids offer a workaround for the issue of non-convexity concerning frequency parameters [55], or alternative metrics are proposed to mitigate the risk of falling into bad local minima [56, 57]. We adopt a parameter regression joint training strategy, akin to the pre-training phase of the work by Engel et al. [58]. Among the output mode frequencies, we train the model to match one mode frequency component (denoted by $\hat{f}_{0}$) to match the fundamental frequency of the target FDTD-simulated audio ($f_{0}$) annotated using CREPE [59] as $\mathcal{L}_{f_{0}}=\|\hat{f}_{0}-f_{0}\|_{1}$.
@@ 4.1 Experimental Setup
Baselines. Table 2 compares the major differences between the baselines and the proposed models. As the first attempt to tackle the task of neural audio synthesis for dynamic physical properties, we compare our proposed model to three other models. We consider three baselines, namely: Modal and DDSPish, and DDSPish-XFM. Modal synthesis is the linear wave solution as in Equation 1, where the modal frequencies and the shape functions are pre-computed. DDSPish is a neural network based on the harmonic plus noise model, similar to DDSP. Yet, this -ish suffix emphasizes that this model is different from the DDSP model [16]. DDSPish does not have a reverb module but instead adds frequency modulation, and most notably it has a parameter encoder that allows generating sounds from physical parameters. Please see Appendix D for more details of the baselines.
Evaluation Metrics. We report results on three metrics, signal-distortion-ratio (SDR), scale-invariant signal-distortion-ratio (SI-SDR) [60], multi-scale spectral (MSS) distance, and the pitch difference in Hz. As with the problem statement, we consider the FDTD-simulated results as the ground truth (GT). Both SDR and SI-SDR estimate the distortions on a spatiotemporal grid and are very strict about scoring out-of-phase cases, by directly comparing the the estimated displacements to the FDTD results without any specific transformations or interpolations. Given that the magnitudes of the estimated displacements typically distribute within the small range (approximately between ±$\pm 0.02$; depends on the pluck amplitude $p_{a}$, see Table 5), we compare both with scale normalization (SI-SDR) and without (SDR). The MSS metric for the evaluation is computed using the short-time Fourier transformation (STFT) with three scales of Fast Fourier Transform (FFT) points: 1024, 512, and 256, with a window length equal to the FFT point for each, and a hop length equal to a quarter of that. For each scale, the STFT magnitude is compared on both linear and log scales, weighted by 2.0 and 0.5, respectively. This choice of weighting is to help scale the loss computation with MSS and should not have a significant impact on performance, but is important to note for absolute comparisons of MSS scores. The Pitch metric is computed as the $\ell^{1}$ norm of the difference between $\hat{f}_{0}$ and $f_{0}$.
Table 2: Comparison between the baselines and the proposed models.
Method Center frequency Network architecture Training configuration
AM block FM block $\mathcal{L}_{1}$ MSS $\mathcal{L}_{f_{0}}$ $\mathcal{L}_{m}$
Modal Mode frequency N/A N/A N/A N/A
DDSPish-XFM Integer multiples
DDSPish Integer multiples
DMSP-Hybrid Mode frequency
DMSP Learnt estimates
@@ 4.2 Results
Differentiable Sound Synthesis. The efficacy of DMSP is studied as shown in Table 3. First, the modal synthesis method calculates a linear solution for a damped stiff string, and the mode for the linear test set is calculated offline. The discrepancy between the outcomes of modal synthesis and FDTD in the linear case can be attributed to the absence of frequency-dependent damping in Equation 1. In other words, the frequency information serves as an oracle for the Modal, as evidenced by its lowest pitch score, but the decay of amplitude over time is not sufficiently modeled, which is where the remaining models demonstrate superior performance. Considering the DDSPish models, the MSS is approximately 1.6 dB ahead of the DMSPs, but the difference can reach up to 44 dB in the case of the SI-SDR. It is worth noting that $\alpha$ is uniformly sampled within the range $(1,25)$ for the training data, falling into the category of the nonlinear strings. Information about the ranges for sampling PDE parameters is in Table 5. In the case of the nonlinear test set, the difference between the modal solution and the FDTD solution becomes larger. On the other hand, DDSPish models, which are based on spectral modeling synthesis but learned without using any modal information, show the lowest performance. The superiority of DMSP is even more pronounced in the nonlinear case. While Modal, which can only cover the solution to the linear case, shows attenuated performance, the DMSPs demonstrate the best performance in all metrics. In particular, DMSP-Hybrid, which precomputes the mode frequency like Modal, performs FM and AM for the nonlinear solution, showing an SI-SDR improvement of nearly 31 dB and an MSS improvement of nearly 13 dB over Modal. DMSP, which estimates the mode as a neural network without the pre-computation step, also outperforms Modal on all metrics for the nonlinear case. The primary reason for the performance discrepancy between the nonlinear case and the linear case for DMSP is that the training data is rarely precisely equal to $1$ in the $\alpha$ distribution when it is sampled.
@@
Figure 5: Objective scores over the change of physical parameters.
@@
Controllable Physical Simulation. The quantitative scores for various physical condition parameters are visualized in Figure 5. Trends show how the results of Modal synthesis and DMSP vary for different pickup positions ($x$), stiffness ($\kappa$), tension ($\alpha$), pluck amplitude ($p_{a}$), and pluck position ($p_{x}$). Of these, $\alpha$ and $p_{a}$ in particular are known to increase the nonlinearity of the string as they increase in magnitude, which can be seen by the lower Modal synthesis scores. For DMSP, we see an overall improvement in the score, with a lower propensity for nonlinearity. Figure 4 depicts the simulated state of the string as the pluck position in the initial condition is varied. The results synthesized by DMSP can reconstruct a very accurate initial condition, similar to the results simulated by FDTD. The vibration propagating through time along the string exhibits a distinct behavior contingent upon the initial condition. FDTD employs a recursive calculation of displacement, necessitating some iterations equal to the number of samples at the audio sampling rate. In contrast, DMSP is capable of obtaining the desired displacement in both time and space simultaneously.
@@
String Sound Synthesis. Spectrograms of the test samples are visualized in Figure 6 and Figure 7. For the spectrograms, the instantaneous frequencies are identified in a rainbow color map, where the color intensities represent the logarithmic magnitude of the power spectra. Observing from Figure 6, the FDTD-simulated spectrogram clearly shows pitch glide and phantom partials at the beginning of the pluck. In contrast, modal synthesis methods that model linear solutions do not show these nonlinear characteristics. The DDSPish-XFM model employs a harmonic pitch skeleton comprising integer multiples of $f_{0}$, thereby precluding the inharmoicities resulting from stiffness. The DDSPish model demonstrates enhanced mode estimation capabilities through learned FM, which modulates the harmonic pitch skeleton to be inharmonic. However, there is scope for further improvement in frequency estimation, particularly in instances where the FM learning process is unstable for high frequencies. On the other hand, the DMSP model, which estimates the mode frequency and amplitude from $u_{0}$, shows an improved pitch skeleton and stable frequency estimation. The DMSP-Hybrid model, which learns to AM and FM the sinusoidal oscillators of the modal solution that requires mode precomputation, shows the most similar results to FDTD.
Figure 7 shows a similar trend for the spectrograms. In this example, however, we reveal a key difference that is not apparent in the spectrograms, but is very important for physical modeling: It is the measure of displacement to space, namely the state plot. The timestep for the state plot can be arbitrary, but we show the initial state for ease of comparison. For methods that do not have a separate method for accurately predicting the mode information, unlike DMSP, difficulties can be found in accurately predicting the state. This partially explains the SI-SDR scores for DDSPish and DDSPish-XFM in Table 3. Considering that the major difference between DMSP and DDSPish is whether the input frequency of the FM block is the mode frequency or an integer multiple, it can be inferred that accurately estimating the mode frequency is critical from a physical modeling perspective to fit the displacement of the strings over time.
@@
(a) Modal
(b) DDSPish-XFM
(c) DDSPish
(d) DMSP
(e) DMSP-Hybrid
(f) FDTD (GT)
Figure 7: Spectrograms and state samples of the synthesized samples on the test set. For the spectrograms shown in the first column, the intensity of the frequency (vertical axis) component for time (horizontal axis) is expressed as brightness, and for the states shown in the second column, the displacement (vertical axis) for space (horizontal axis).
@@
Ablation Study. The ablation study on the choice of the training loss functions is presented in Table 4. Overall, for linear strings, the performance does not vary much depending on which loss is used, especially for pitch. This is due to a gating applied to the FM block, which is designed to prevent FM from occurring when $\alpha$ is 1. More specifically, we apply a gating, e.g., $\tanh(\alpha-1)$, in such a way that frequency modulation only occurs when the value of $\alpha$ deviates from 1, which can directly affect pitch, and otherwise forces the mode estimation result to be used as is. This is the main factor that determines the nonlinearity of the string. For AM, there is no such masking depending on $\alpha$, so the remaining metrics except pitch do vary. For nonlinear data, the training results vary depending on the loss design. In particular, for all metrics, using $\mathcal{L}_{f_{0}}$ loss significantly improved performance over not using it. This difference is especially noticeable in the MSS scores between DMSP-Hybrid w.o. $\mathcal{L}_{f_{0}}$ and DMSP, where DMSP even gives better results over the model that uses precomputed mode frequencies for nonlinear MSS scores if $\mathcal{L}_{f_{0}}$ is applied. This result reaffirms the validity of the FM block for nonlinear strings in terms of predicting dynamically varying frequencies such as pitch glide and shows that $\mathcal{L}_{f_{0}}$ loss is the loss function to train it effectively.
@@
Motion Synthesis. The main advantage of DMSP, compared to the existing sound synthesis models, is that it can synthesize not only sound but also motion, which is one of the main characteristics of physical modeling techniques. In particular, the DMSP can visualize the corresponding string motion as a video, although the receptive field and computational complexity required to obtain a solution $u(x,t)$ for a single spatio-temporal point is of order 1, as shown in Table 1, when these solutions are pooled for a given $x\in\Omega$ and $t\in[0,1)$, the corresponding string motion can be visualized as a video. Figure 8 visualizes the resulting transverse displacement of the string over time (horizontal axis) and space (vertical axis). The transverse displacement of the FDTD is coupled to the longitudinal motion, which is why it differs from the Modal synthesis output, which synthesizes the motion of a linear damped stiff string. The results output by DMSP show improved accuracy.